# Negate or Embrace: On How Misalignment Shapes Multimodal Representation Learning

### v1

@@ Results.
Figure 6 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably captured across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$. Likewise, semantic variables omitted due to selection or perturbation bias are generally discarded. For instance, in selection setting 1 or perturbation setting 7, only shape remains predictable. When factors such as x_pos or y_pos are included, other dependent semantic variables, such as color, become partially predictable (e.g., in selection setting 2). Similarly, identification of s_pos enhances predictability of s_color and b_color, reflecting the latent causal structure illustrated in Figure 14. Overall, the results on the Causal3DIdent dataset further support our theoretical findings. Analyses of the text representations are provided in § F.2.
@@ 6 Conclusion and Discussion
In this work, we present a formal analysis of misalignment in MMCL, examining its impact on learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, systematically discarding misaligned latent variables. When image-text pairs exhibit selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights underscore the need for multimodal learning frameworks that either mitigate misalignment or leverage beneficial biases to improve representation learning in real-world settings (see App. H for further discussion).
@@ Appendix A Notation and Terminology
$\mathbf{g}_{x}$ Generative mapping for images: $\mathcal{S}\times\mathcal{M}_{x}\to\mathcal{X}$
$\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ Generative mapping for text under selection bias $\theta$: $\mathcal{S}_{\mathbb{I}_{\theta}}\times\mathcal{M}_{t}\to\mathcal{T}^{\scriptscriptstyle(\theta)}$
@@
$\mathcal{L}_{\text{\tiny MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq. 1)
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq. 2)
@@
Appendix B Related Work
Theoretical Multimodal (Multi-View) Contrastive Learning.
Recent work has sought to formalize the theoretical foundations of multimodal and multi-view representation learning, particularly under contrastive objectives. Wang et al. [67] decompose the InfoNCE loss [49] into an alignment term, which pulls positive pairs together, and a uniformity term, which encourages dispersion over a hypersphere—laying the groundwork for subsequent analysis. Zimmermann et al. [80] show that contrastive objectives can invert the data-generating process, while Liu et al. [43] extend this result to multimodal settings. However, these approaches often rely on strong assumptions about latent distributions or manifold structures, which limits their practical applicability. A complementary line of work, such as [66], demonstrates that contrastive learning with data augmentation can recover shared content without requiring such strong assumptions. This has been extended to MMCL by [46, 12], and Yao et al. [76] further investigate identifiability under partial observability in multi-view settings. Distinct from prior work, we do not assume fixed content–style decompositions, causal directions, or perfectly aligned pairs. Instead, we analyze MMCL with image–text pairs under misalignment and systematically examine its impact on representation learning.
Vision-Language Models and Perspectives on Misalignment.
Multimodal contrastive learning (MMCL) has achieved significant empirical success, particularly in aligning visual and textual modalities using models such as CLIP [53] and ALIGN [25]. These successes are partly attributed to the use of massive training corpora, e.g., LAION-5B [58], which are substantially larger than those used for vision-only foundation models [21, 50]. However, real-world multimodal datasets are often imperfectly aligned and noisy [47]. Existing empirical methods [1, 32] typically treat such misalignment as label noise, employing strategies such as multiple-instance learning or dataset refinement [1] to mitigate its impact. While some recent work suggests that contrastive models are robust to certain forms of structured misalignment [48], others propose augmenting text to simulate semantic variation in visual content [5]. Our work suggests that misalignment can act as either a barrier or a bridge, depending on the application. Unlike [48], which assumes linear representations without modeling the generative process, our analysis is grounded in a realistic latent variable model that provides a deeper understanding of misalignment from a data-generating perspective.
Identifiability in Latent Variable Models.
Identifiability analysis addresses the fundamental question of whether the learning process can uniquely recover the latent generative structure or distribution underlying the observed data. This problem has been extensively studied in the context of nonlinear independent component analysis (ICA) [22, 44, 26, 60] and causal representation learning [39, 41, 77, 75, 42]. In practice, full identifiability—typically up to permutation—is rarely achievable without strong assumptions. Consequently, recent works have focused on partial identifiability [19, 29, 17] or relaxed equivalence classes, such as identifiability up to linear transformations [80, 43] or up to group-wise/block-wise indeterminacy [66, 76], which can offer sufficient guarantees for specific tasks or settings. In the context of multimodal representation learning, several recent studies have explored identifiability results [46, 12, 43], but largely neglect the presence of systematic misalignment. In contrast, our work explicitly models misalignment and adopts a block-identifiability definition to characterize the extent to which semantic factors can be recovered up to an invertible mapping.
Invariant Representation Learning.
Invariant representation learning (IRL) seeks to learn representations that remain robust under distributional shifts between environments [2, 14, 17], particularly in settings where empirical risk minimization (ERM) [64] fails to generalize out of distribution. In the absence of such variation, ERM is sufficient for in-distribution prediction, rendering the objective of IRL ill-posed. From a causal perspective, learning invariant representations—or more ambitiously, invariant mechanisms—requires variability in the non-invariant factors of the data. Such variability can be introduced through interventional data [37, 31], exchangeability assumptions [54], or the use of auxiliary variables such as domain indices [77, 40]. However, direct interventions on latent variables are typically infeasible in real-world data, and auxiliary variable methods often require access to a large number of diverse environments to ensure identifiability—an assumption that is rarely satisfied in practice. Our work offers an alternative approach: by leveraging the inherent flexibility of text supervision, we demonstrate that manipulating biases—specifically through selective omission or semantic perturbation in text—can serve as a controllable proxy for environmental variation.
Appendix C Proofs
C.1 Lemmas
@@
Lemma C.1 (Global Minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$).
@@
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}\Bigr{)},$ (6)
@@
  • $\mathbf{d}=(d_{1},\dots,d_{n})$ is defined via the Darmois construction [11, 22, 66], where for each $i\in[n]$ (we abbreviate $\{1,\cdots,n\}$ as $[n]$ for any integer $n>1$ for simplicity),
@@
Proof of Lem. C.1.
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations 7 and 8 yield $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq. 6, we have
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\Bigl{[}\|\mathbf{f}_{x}^{*}(\mathbf{x})-\mathbf{f}_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\Bigr{]}$
@@
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}}}\Bigl{[}\|\mathbf{d}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}})-\mathbf{d}(\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}})\|_{2}\Bigr{]}$
@@
Further, by the properties of the Darmois construction [11], the mapping $\mathbf{d}$ transforms $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ into a uniform distribution over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$). Since the uniform distribution is the unique maximum entropy (i.e., zero) distribution on a bounded domain (under no further moment constraints) [24, 9], the entropy terms in the loss are maximized. In the formulation of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$, this maximal entropy precisely cancels any potential reduction in the loss, ensuring that
$\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0.$
Therefore, the global minimum of $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ is achieved at 0 by the given function pairs $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$, completing the proof. ∎
Lemma C.2 (Uniformizing Mapping Preserves All Information).
@@
Proof of Lem. C.2.
Let $p_{\mathbf{u}}:\mathcal{U}\to\mathbb{R}$ and $p_{\mathbf{v}}:\mathcal{V}\to\mathbb{R}$ denote the probability density functions of $\mathbf{u}$ and $\mathbf{v}$, respectively. Since $\mathbf{v}=\mathbf{h}(\mathbf{u})$, the change-of-variables formula (refer to, e.g., 4) yields
@@
or equivalently, $\det J(\mathbf{h})(\mathbf{u})\neq 0$ a.e.. By the Inverse Function Theorem (see, e.g., 56), this implies that $\mathbf{h}$ is a local diffeomorphism.
Moreover, since $\mathcal{U}$ and $\mathcal{V}$ are simply connected, open $\mathcal{C}^{1}$ manifolds, standard covering space theory (refer to, e.g., the discussion around Theorem 1.38 in 20) implies that $\mathbf{h}$ is a covering map. The uniformity of $p_{\mathbf{v}}$ forces $\mathbf{h}$ to be surjective (otherwise, some points in $\mathcal{V}$ would have zero density, contradicting uniformity). Since any covering map from a simply connected space is trivial, $\mathbf{h}$ must be equivalent to the identity covering. In other words, $\mathbf{h}$ is a homeomorphism onto $\mathcal{V}$ and hence both injective and surjective (i.e., a bijection).
Finally, the fact that the Jacobian determinant is nonzero a.e. guarantees that $\mathbf{h}$ depends on all components of $\mathbf{u}$; if any component were omitted, the rank of the Jacobian would drop, contradicting non-singularity. Furthermore, by the Global Inverse Function Theorem (refer to, e.g., 57), the inverse of $\mathbf{h}$ is smooth.
@@
    First, we show that the objective function $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 00. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
    Next, we prove that minimizing $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ inherently eliminates any dependence of the learned representations on modality-specific variables $\mathbf{m}_{x}$ or $\mathbf{m}_{t}$. This ensures that the representations are restricted to the dependence on latent semantic variables.
@@
Since all terms in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ are non-negative, and its global minimum is 00 by Lem. C.1, each term in $\mathcal{L}_{\text{\tiny SymAlignMaxEnt}}$ must vanish a.s. for any pairing $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, leading to:
@@
Clarification C.1 (Causal Interpretations). (i) Justification for the existence of distinct points in Eq. 20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm. 3.1. From a latent SCM perspective [51, 65], even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
@@
By directly applying Lem. C.2, it follows that the learned representations $\hat{\mathbf{z}}_{x}$ (and also $\hat{\mathbf{z}}_{t}$) include all and only the information of the unaltered semantic components $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ almost surely, and that $\mathbf{h}_{x}$ (and similarly $\mathbf{h}_{t}$) is invertible. Consequently, the true modality-shared semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ are block-identified by $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
@@
Since both the omitted and perturbable index subsets are trivial, we may directly apply Lem. C.2, which implies that the full semantic vector $\mathbf{s}$ is block-identified by both $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.

### v2

@@ Results.
Figure 6 presents the prediction performance of a nonlinear MLP classifier or regressor trained on learned image representations. We observe that unbiased semantic variables, whether continuous (e.g., s_pos) or discrete (e.g., shape, x_pos, y_pos), are reliably captured across all settings, with predictive performance approaching perfect ($R^{2}\approx 1$). For semantic variables that are continuous in the image modality but discretized in the text modality, prediction performance shows some degradation (e.g., s_color in selection setting 6 or perturbation setting 2), yet still achieves relatively high $R^{2}$ scores. Image-specific variables are consistently excluded from the learned representations, as indicated by $R^{2}=0$. Likewise, semantic variables omitted due to selection or perturbation bias are generally discarded. For instance, in selection setting 1 or perturbation setting 7, only shape remains predictable. When factors such as x_pos or y_pos are included, other dependent semantic variables, such as color, become partially predictable (e.g., in selection setting 2). Similarly, identification of s_pos enhances predictability of s_color and b_color, reflecting the latent causal structure illustrated in Figure 14. Overall, the results on the Causal3DIdent dataset further support our theoretical findings. Analyses of the text representations are provided in § E.2.
@@
6 Related Work
Theoretical Multimodal (Multi-View) Contrastive Learning.
Recent work has sought to formalize the theoretical foundations of multimodal and multi-view representation learning, particularly under contrastive objectives. Wang et al. (Wang and Isola, 2020) decompose the InfoNCE loss (Oord et al., 2018) into an alignment term, which pulls positive pairs together, and a uniformity term, which encourages dispersion over a hypersphere—laying the groundwork for subsequent analysis. Zimmermann et al. (Zimmermann et al., 2021) show that contrastive objectives can invert the data-generating process, while Liu et al. (Liu et al., 2024d) extend this result to multimodal settings. However, these approaches often rely on strong assumptions about latent distributions or manifold structures, which limits their practical applicability. A complementary line of work, such as (Von Kügelgen et al., 2021), demonstrates that contrastive learning with data augmentation can recover shared content without requiring such strong assumptions. This has been extended to MMCL by (Lyu et al., 2022; Daunhawer et al., 2022), and Yao et al. (Yao et al., 2023) further investigate identifiability under partial observability in multi-view settings. Distinct from prior work, we do not assume fixed content–style decompositions, causal directions, or perfectly aligned pairs. Instead, we analyze MMCL with image–text pairs under misalignment and systematically examine its impact on representation learning.
Vision-Language Models and Perspectives on Misalignment.
Multimodal contrastive learning (MMCL) has achieved significant empirical success, particularly in aligning visual and textual modalities using models such as CLIP (Radford et al., 2021) and ALIGN (Jia et al., 2021). These successes are partly attributed to the use of massive training corpora, e.g., LAION-5B (Schuhmann et al., 2022), which are substantially larger than those used for vision-only foundation models (He et al., 2022; Oquab et al., 2024). However, real-world multimodal datasets are often imperfectly aligned and noisy (Miech et al., 2019). Existing empirical methods (Alayrac et al., 2020; Lavoie et al., 2024) typically treat such misalignment as label noise, employing strategies such as multiple-instance learning or dataset refinement (Alayrac et al., 2020) to mitigate its impact. While some recent work suggests that contrastive models are robust to certain forms of structured misalignment (Nakada et al., 2023), others propose augmenting text to simulate semantic variation in visual content (Cai et al., 2025). Our work suggests that misalignment can act as either a barrier or a bridge, depending on the application. Unlike (Nakada et al., 2023), which assumes linear representations without modeling the generative process, our analysis is grounded in a realistic latent variable model that provides a deeper understanding of misalignment from a data-generating perspective.
Identifiability in Latent Variable Models.
Identifiability analysis addresses the fundamental question of whether the learning process can uniquely recover the latent generative structure or distribution underlying the observed data. This problem has been extensively studied in the context of nonlinear independent component analysis (ICA) (Hyvärinen and Pajunen, 1999; Locatello et al., 2019; Khemakhem et al., 2020; Sorrenson et al., 2020) and causal representation learning (Liu et al., 2022, 2024c; Zhang et al., 2024a; Yao et al., 2025; Liu et al., 2025). In practice, full identifiability—typically up to permutation—is rarely achievable without strong assumptions. Consequently, recent works have focused on partial identifiability (Gu and Xu, 2020; Kong et al., 2023; Gao et al., 2025) or relaxed equivalence classes, such as identifiability up to linear transformations (Zimmermann et al., 2021; Liu et al., 2024d) or up to group-wise/block-wise indeterminacy (Von Kügelgen et al., 2021; Yao et al., 2023), which can offer sufficient guarantees for specific tasks or settings. In the context of multimodal representation learning, several recent studies have explored identifiability results (Lyu et al., 2022; Daunhawer et al., 2022; Liu et al., 2024d), but largely neglect the presence of systematic misalignment. In contrast, our work explicitly models misalignment and adopts a block-identifiability definition to characterize the extent to which semantic factors can be recovered up to an invertible mapping.
Invariant Representation Learning.
Invariant representation learning (IRL) seeks to learn representations that remain robust under distributional shifts between environments (Arjovsky et al., 2019; Eastwood et al., 2022; Gao et al., 2025), particularly in settings where empirical risk minimization (ERM) (Vapnik, 1991) fails to generalize out of distribution. In the absence of such variation, ERM is sufficient for in-distribution prediction, rendering the objective of IRL ill-posed. From a causal perspective, learning invariant representations—or more ambitiously, invariant mechanisms—requires variability in the non-invariant factors of the data. Such variability can be introduced through interventional data (Lippe et al., 2022; Lachapelle et al., 2022), exchangeability assumptions (Reizinger et al., 2025), or the use of auxiliary variables such as domain indices (Zhang et al., 2024a; Liu et al., 2024b). However, direct interventions on latent variables are typically infeasible in real-world data, and auxiliary variable methods often require access to a large number of diverse environments to ensure identifiability—an assumption that is rarely satisfied in practice. Our work offers an alternative approach: by leveraging the inherent flexibility of text supervision, we demonstrate that manipulating biases—specifically through selective omission or semantic perturbation in text—can serve as a controllable proxy for environmental variation.
@@ 7 Conclusion and Discussion
In this work, we present a formal analysis of misalignment in MMCL, examining its impact on learned representations. We demonstrate that contrastive multimodal encoders retain only the unbiased shared semantic variables, systematically discarding misaligned latent variables. When image-text pairs exhibit selection or perturbation biases, the joint embedding prioritizes consistent content, while omitting altered or missing aspects. This trade-off is fundamental: perfectly aligned text captions preserve rich semantic detail, whereas selective or biased text can enhance domain invariance by filtering out distribution-sensitive factors. Our experiments, conducted across simulations and image-text datasets, empirically validate these theoretical findings. These insights underscore the need for multimodal learning frameworks that either mitigate misalignment or leverage beneficial biases to improve representation learning in real-world settings (see App. G for further discussion).
@@ Appendix A Notation and Terminology
$\mathbf{g}_{x}$ Differomorphic generative mapping for images: $\mathcal{S}\times\mathcal{M}_{x}\to\mathcal{X}$
$\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ Differomorphic generative mapping for text under selection $\theta$: $\mathcal{S}_{\mathbb{I}_{\theta}}\times\mathcal{M}_{t}\to\mathcal{T}^{\scriptscriptstyle(\theta)}$
$\mathcal{G}_{t}$ A diffeomorphism function class where $\mathbf{g}_{t^{\scriptscriptstyle(\theta)}}$ is selected from
@@
$\mathcal{L}_{\text{MMCL}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Symmetric InfoNCE loss for MMCL (Eq. 1)
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ Alignment and entropy maximization loss (Eq. 2)
@@
Appendix B Proofs
B.1 Lemmas
@@
Lemma B.1 (Global Minimum of $\mathcal{L}_{\text{SymAlignMaxEnt}}$).
@@
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\left[\|\mathbf{f}_{x}(\mathbf{x})-\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\right]-\frac{1}{2}\Bigl{(}H\bigl{(}\mathbf{f}_{x}(\mathbf{x})\bigr{)}+H\bigl{(}\mathbf{f}_{t}(\mathbf{t}^{\scriptscriptstyle(\theta)})\bigr{)}\Bigr{)},$ (6)
@@
  • $\mathbf{d}=(d_{1},\dots,d_{n})$ is defined via the Darmois construction (Darmois, 1951; Hyvärinen and Pajunen, 1999; Von Kügelgen et al., 2021), where for each $i\in[n]$ (we abbreviate $\{1,\cdots,n\}$ as $[n]$ for any integer $n>1$ for simplicity),
@@
Proof of Lem. B.1.
We prove that the candidate functions $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$ in Equations 7 and 8 yield $\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0$. Substituting these candidate functions into the loss in Eq. 6, we have
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})\sim p_{\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)}}}\Bigl{[}\|\mathbf{f}_{x}^{*}(\mathbf{x})-\mathbf{f}_{t}^{*}(\mathbf{t}^{\scriptscriptstyle(\theta)})\|_{2}\Bigr{]}$
@@
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})$ $=\mathop{\mathbb{E}}_{(\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}})\sim p_{\mathbf{z}_{x},\mathbf{z}_{t^{\scriptscriptstyle(\theta)}}}}\Bigl{[}\|\mathbf{d}(\mathbf{s}_{\mathbb{I}^{c}_{\rho}})-\mathbf{d}(\tilde{\mathbf{s}}_{\mathbb{I}^{c}_{\rho}})\|_{2}\Bigr{]}$
@@
Further, by the properties of the Darmois construction (Darmois, 1951), the mapping $\mathbf{d}$ transforms $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ into a uniform distribution over $(0,1)^{n}$ (with $n=|\mathbb{I}^{c}_{\rho}|$). Since the uniform distribution is the unique maximum entropy (i.e., zero) distribution on a bounded domain (under no further moment constraints) (Jaynes, 1982; Cover, 2006), the entropy terms in the loss are maximized. In the formulation of $\mathcal{L}_{\text{SymAlignMaxEnt}}$, this maximal entropy precisely cancels any potential reduction in the loss, ensuring that
$\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x}^{*},\mathbf{f}_{t}^{*})=0.$
Therefore, the global minimum of $\mathcal{L}_{\text{SymAlignMaxEnt}}$ is achieved at 0 by the given function pairs $\mathbf{f}_{x}^{*}$ and $\mathbf{f}_{t}^{*}$, completing the proof. ∎
Lemma B.2 (Uniformizing Mapping Preserves All Information).
@@
Proof of Lem. B.2.
Let $p_{\mathbf{u}}:\mathcal{U}\to\mathbb{R}$ and $p_{\mathbf{v}}:\mathcal{V}\to\mathbb{R}$ denote the probability density functions of $\mathbf{u}$ and $\mathbf{v}$, respectively. Since $\mathbf{v}=\mathbf{h}(\mathbf{u})$, the change-of-variables formula (refer to, e.g., Bogachev, 2007) yields
@@
or equivalently, $\det J(\mathbf{h})(\mathbf{u})\neq 0$ a.e.. By the Inverse Function Theorem (see, e.g., Rudin, 1976), this implies that $\mathbf{h}$ is a local diffeomorphism.
Moreover, since $\mathcal{U}$ and $\mathcal{V}$ are simply connected, open $\mathcal{C}^{1}$ manifolds, standard covering space theory (refer to, e.g., the discussion around Theorem 1.38 in Hatcher, 2002) implies that $\mathbf{h}$ is a covering map. The uniformity of $p_{\mathbf{v}}$ forces $\mathbf{h}$ to be surjective (otherwise, some points in $\mathcal{V}$ would have zero density, contradicting uniformity). Since any covering map from a simply connected space is trivial, $\mathbf{h}$ must be equivalent to the identity covering. In other words, $\mathbf{h}$ is a homeomorphism onto $\mathcal{V}$ and hence both injective and surjective (i.e., a bijection).
Finally, the fact that the Jacobian determinant is nonzero a.e. guarantees that $\mathbf{h}$ depends on all components of $\mathbf{u}$; if any component were omitted, the rank of the Jacobian would drop, contradicting non-singularity. Furthermore, by the Global Inverse Function Theorem (refer to, e.g., Ruzhansky and Sugimoto, 2015), the inverse of $\mathbf{h}$ is smooth.
@@
    First, we show that the objective function $\mathcal{L}_{\text{SymAlignMaxEnt}}(\mathbf{f}_{x},\mathbf{f}_{t})$ achieves a global minimum value of 00. At this minimum, any pair of smooth functions $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$ satisfying this condition must exhibit invariance across modalities. This invariance condition ensures that the learned image representations and text representations must align across all positive $\mathbf{x}$ and $\mathbf{t}^{\scriptscriptstyle(\theta)}$ pairs.
@@
    Next, we prove that minimizing $\mathcal{L}_{\text{SymAlignMaxEnt}}$ inherently eliminates any dependence of the learned representations on modality-specific variables $\mathbf{m}_{x}$ or $\mathbf{m}_{t}$. This ensures that the representations are restricted to the dependence on latent semantic variables.
@@
Since all terms in $\mathcal{L}_{\text{SymAlignMaxEnt}}$ are non-negative, and its global minimum is 00 by Lem. B.1, each term in $\mathcal{L}_{\text{SymAlignMaxEnt}}$ must vanish a.s. for any pairing $(\mathbf{x},\mathbf{t}^{\scriptscriptstyle(\theta)})$, leading to:
@@
Clarification B.1 (Causal Interpretations).
(i) Justification for the existence of distinct points in Eq. 20. This follows from the assumption $p_{\mathbf{s}}>0$ a.e. on $\mathcal{S}$ by Asm. 3.1. From a latent SCM perspective (Pearl, 2009; von Kügelgen et al., 2024), even if a specific semantic component $s_{l}$ in $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ is the ancestor node of some other semantic components in $\mathbf{s}_{\mathbb{I}_{\theta}}$, the strict positivity of $p_{\mathbf{s}}$ ensures that the exogenous noise variables are well-defined. Thus, for different values of $s_{l}$, there exist corresponding noise values that keep $\mathbf{s}^{*}_{\mathbb{I}_{\theta}}$ remaining fixed. (ii) What if the unknown causal structure is $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}\to\mathbf{s}_{\mathbb{I}_{\theta}}$? The potential causal influence from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to $\mathbf{s}_{\mathbb{I}_{\theta}}$ does not resolve the contradiction. Independence here means that, once $\mathbf{s}_{\mathbb{I}_{\theta}}$ is set, there is no direct functional path from $\mathbf{s}_{\mathbb{I}^{c}_{\theta}}$ to the representations $\mathbf{h}_{x}(\mathbf{s}_{\mathbb{I}_{\theta}})$, i.e., the causal influence among them is fully accounted for by the realized value of $\mathbf{s}_{\mathbb{I}_{\theta}}$.
@@
By directly applying Lem. B.2, it follows that the learned representations $\hat{\mathbf{z}}_{x}$ (and also $\hat{\mathbf{z}}_{t}$) include all and only the information of the unaltered semantic components $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ almost surely, and that $\mathbf{h}_{x}$ (and similarly $\mathbf{h}_{t}$) is invertible. Consequently, the true modality-shared semantic variables $\mathbf{s}_{\mathbb{I}^{c}_{\rho}}$ are block-identified by $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.
@@
Since both the omitted and perturbable index subsets are trivial, we may directly apply Lem. B.2, which implies that the full semantic vector $\mathbf{s}$ is block-identified by both $\mathbf{f}_{x}$ and $\mathbf{f}_{t}$.